# Information theory for data-driven model reduction in physics and biology

### v2

@@ Appendix I Cyanobacteria experiments
$r(t)\text{e}^{i\psi(t)}=\frac{1}{V}\left|\int\text{d}\mathbf{x}\text{e}^{i\theta(\mathbf{x},t)}\right|$ (83)
@@
Appendix J Simulation Parameters
Triple well simulations For the dynamics of the triple well we work directly with the force $F=-\partial_{x}U$,
@@
$\text{d}x_{t}=\frac{F(x_{t})}{\gamma}\text{d}t+\sigma\sqrt{\text{d}t}\eta_{t}$ (84)
where $\eta_{t}$ is white noise with unit variance. The noise magnitude $\sigma$ is related to the diffusion constant in Eq. 26 by $D=\frac{\sigma^{2}}{2}$. We use $\sigma=1.0$ and $\gamma=0.2$
@@
Pitchfork bifurcation simulations The system was evolved according to the stochastic differential equation Eq. 84 with $F(x)=-\mu x-x^{3}$, $\gamma=1$ and $\sigma=0.1$. For each value of $\mu$, $10^{5}$ initial conditions were simulated with 2000 trajectories starting at each initial condition. These were evolved for 100 time steps of $\text{d}t=2\cdot 10^{-3}$.
Hopf oscillator simulations The system was evolved according to Eq. 84 with the force given by Eq.68 in the main text, with $\gamma=1$ and $\sigma=10^{-2}$, $\omega=4.0$ and $\mu=0.25$. For each value of $\mu$, $10^{6}$ initial conditions were simulated with 1000 trajectories starting at each initial condition. These were evolved for 50 time steps of $\text{d}t=2\cdot 10^{-3}$.
Lorenz system simulations The system was evolved according to
$\dot{x}$ $=\sigma(y-x)$ (85)
$\dot{y}$ $=x(\rho-z)-y$ (86)
$\dot{z}$ $=xy-\beta z$ (87)
@@
Fluid flow simulations The fluid flow simulations simulations are contained in the “Cylinder in Crossflow” dataset, downloaded from Ref. [93]. We take the dataset at Reynolds number 150, and interpolate the velocity field from the unstructured 6569-node mesh to a regular grid of $300\times 150$ pixels. The VIB networks are trained with $\beta=10^{7}$ and a time buffer $\Delta t$ chosen randomly between $[4,24]$ (see our comment on time randomization in “VIB for deterministic dynamics” in Section F).
Locally-coupled Kuramoto model In SI Fig. S11 we perform variational IB on a locally-coupled Kuramoto model and find the latent variables learn a synchronization order parameter as in the cyanobacteria data. This model considers a spatially-distributed field of phases which evolves according to
@@
Appendix K Comparison of various dimensionality reduction methods
In this section we compare the performance of variational IB (VIB) to several common data-driven model reduction or inference methods. VIB uses a neural network to identify the relevant variables that are most predictive of the future. This construction learns dynamically relevant variables in contrast to methods such as principal component analysis or diffusion maps (in the absence of additional stringent assumptions), and it finds potentially non-linear variables in an agnostic way without requiring a tailored library of non-linearities (such as extended dynamic mode decomposition). In addition to these factors, VIB produces physically well-defined relevant variables – transfer operator eigenfunctions – in contrast to other deep methods such as (variational) autoencoders. In Fig. S12 we compare the low-dimensional latent space trajectories of the reduced models obtained using several of these standard methods to those output by VIB. In the remainder of this section we discuss these methods in detail.
@@
The principal components of the data are eigenvectors of the covariance matrix; the dominant principal component determines the direction of most variation in the dataset. In practice, if many samples are present the correlation matrix is expensive to compute. Writing the dataset as a matrix, $\mathbf{X}\in\mathbb{R}^{N_{\text{samples}}\times N_{\text{feat}}}$, computing the correlation matrix requires evaluating the matrix product $\mathbf{C}=\mathbf{X}^{T}\mathbf{X}$. Instead, a more computationally efficient way is to compute the singular value decomposition (SVD) of the data matrix,
@@
DMD in its original formulation assumes that the Koopman operator linearly evolves the observed state variable. However, the true Koopman operator evolves arbitrary non-linear functions of the state variable forward in time. DMD has therefore been extended to account for non-linear functions through “extended DMD”, or eDMD [13]. eDMD augments the state vector with non-linear transformations of the state. As an example, a state vector $[-\mathbf{x}-]$ may be replaced by monomials $[-\mathbf{x}-,-\mathbf{x}^{2}-,-\mathbf{x}^{3}-]$ (where here we understand exponentiation as element-wise). Other approaches also exist, such as augmenting the state vector by several time-delayed state vectors [96]. Rather than constructing the full $d\times d$ matrix, where $d$ is the dimension of the (possibly augmented) state, one can directly construct a low-rank approximation of $K$ by computing the reduced-rank singular value decomposition (SVD) of the matrix $K$ from the SVDs of $\mathbf{X}_{t}$ and $\mathbf{X}_{t+\Delta t}$.
@@
$\mathbf{x}=\mathbf{A}_{\text{ICA}}\mathbf{s}.$ (88)
@@
The leveraging of non-Gaussianity in ICA is required due to the fact that, if using whitened data ($\mathbf{z}=\mathbf{V}\mathbf{x}$ so that $\mathbf{z}$ is zero mean ,$\mathbb{E}[\mathbf{z}]=\mathbf{0}$, and unit variance, $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=\mathbf{I}$), then the mixing matrix $\mathbf{A}_{\text{ICA}}$ cannot be estimated if the components $\mathbf{s}$ are normal Gaussian variables: any orthogonal matrix will satisfy (88). For dynamical systems, however, one can get around this requirement of non-Gaussianity. In particular, one can use a time-lagged variable $\mathbf{z}_{t-\tau}=\mathbf{A}_{\text{ICA}}\mathbf{s}_{t-\tau}$ to compute the mixing matrix from the time-correlation matrix of the signal $\mathbf{z}_{t}$ [10]. In particular, we have
@@
where $q_{\phi}(z|x)$ is the posterior on the latent variables $z$ and is parameterized by a neural network (encoder) with parameters $\phi$, and $p_{\theta}(x|z)$ denotes the decoding network with parameters $\theta$. Strictly speaking, we present in the above objective function the $\beta$-VAE loss [83]. With $\beta=1$, which is the case for the original VAE [100], the objective is an upper bound on the log likelihood $-\mathbb{E}[\log p_{\theta}(x)]$ which is minimized when $p_{\theta}(x)$ is equal to the true data distribution $p(x)$. The term $\beta$ controls compression as in the IB objective, however it has the opposite effect: small $\beta_{\text{IB}}$ corresponds to high compression, while small $\beta_{\text{VAE}}$ corresponds to low compression.

### v3

@@ VIII Cyanobacteria experiments
$r(t)\text{e}^{i\psi(t)}=\frac{1}{V}\left|\int\text{d}\mathbf{x}\text{e}^{i\theta(\mathbf{x},t)}\right|$ (S103)
@@
Figure S15: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. The colonies are imaged as they undergo cell growth and oscillations in Kai expression associated with the circadian rhythm. (b) Time series of one cropped section of an individual colony (see SI Fig. S9 for details). (c) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$. (d) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. (e) Embedding into three dimensions. We orient our axes to correspond to the three principal components of the data. Thus, the projection onto the $h_{2}=0$ plane corresponds to a projection along the dominant two principal components. In this subspace, we see a similar structure as in the 2D embedding. (f) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Increasing the dimension of the embedding space beyond two leads only to marginal increases in $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop.” Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. Because the InfoNCE estimator is a lower bound on the true information, we consider the maximum as the estimated mutual information. The value of $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the true dynamics. (g) Selected trajectories in latent space with large and small radii. (h) Time-average latent radius of all colonies. (i) Mean synchronization order parameter of the intensity images; see SI for computation details. (j) Mean radius versus synchronization parameter for each colony. VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. As revealed to us after our analysis, this clustering corresponds to differing theophylline concentrations across experiments. Within each experimental movie (each of which contains 2-3 colonies) the radii are mostly constant. (k) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a slight Gaussian blur to better visualize the bacteria boundaries.
IX Simulation Parameters
Triple well simulations
For the dynamics of the triple well we work directly with the force $F=-\partial_{x}U$,
@@ IX Simulation Parameters
$\text{d}x_{t}=\frac{F(x_{t})}{\gamma}\text{d}t+\sigma\sqrt{\text{d}t}\eta_{t}$ (S104)
where $\eta_{t}$ is white noise with unit variance. The noise magnitude $\sigma$ is related to the diffusion constant in Eq. S10 by $D=\frac{\sigma^{2}}{2}$. We use $\sigma=1.0$ and $\gamma=0.2$
@@
Pitchfork bifurcation simulations
The system was evolved according to the stochastic differential equation Eq. S104 with $F(x)=-\mu x-x^{3}$, $\gamma=1$ and $\sigma=0.1$. For each value of $\mu$, $10^{5}$ initial conditions were simulated with 2000 trajectories starting at each initial condition. These were evolved for 100 time steps of $\text{d}t=2\cdot 10^{-3}$.
Hopf oscillator simulations
The system was evolved according to Eq. S104 with the force given by Eq.S96 above (transformed into cartesian coordinates), with $\gamma=1$ and $\sigma=10^{-2}$, $\omega=4.0$ and $\mu=0.25$. For each value of $\mu$, $10^{6}$ initial conditions were simulated with 1000 trajectories starting at each initial condition. These were evolved for 50 time steps of $\text{d}t=2\cdot 10^{-3}$.
Lorenz system simulations
The system was evolved according to
$\dot{x}$ $=\sigma(y-x)$ (S105)
$\dot{y}$ $=x(\rho-z)-y$ (S106)
$\dot{z}$ $=xy-\beta z$ (S107)
@@
Driven sine-Gordon simulations
The equation of motion of the continuous field $\phi(x)$ is [78]
$\partial_{t}^{2}\phi-\partial_{x}^{2}\phi+\sin\phi$ $=\varepsilon[-\alpha\partial_{t}\phi+\Gamma\cos(\omega t)].$ (S108)
The system is periodic in space, and was simulated with initial conditions
$\phi(x,t)\big|_{t=0}$ $=4\arctan\left[\frac{\sqrt{1-\omega_{\text{br}}^{2}}}{\omega_{\text{br}}\cosh\left(\sqrt{1-\omega_{\text{br}}^{2}}(x-L/2)\right)}\right]$ (S109)
$\partial_{t}\phi(x,t)\big|_{t=0}$ $=0$ (S110)
where $\omega_{\text{br}}=0.77$ is the frequency of the breather initial condition and $L$ is the system size. Key simulation parameters are $\text{d}t=10^{-2}$, $L=24$ (discretized onto $N_{\text{bins}}=256$ bins), $\varepsilon\alpha=0.04$, $\varepsilon\Gamma=0.105$, and $\omega=0.87$ (see Ref. [78]).
Kuramoto-Sivashinsky simulations
The Kuramoto Sivashinsky model describes the evolution of a one-dimensional field $u(x)$ according to
$\partial_{t}u=-\partial_{x}^{4}u-\partial_{x}^{2}u-u\partial_{x}u$ (S111)
Key simulation parameters are the system size $L=22$, time resolution $\text{d}t=2\cdot 10^{-3}$, total simulation time $T=300$, and spatial resolution $N_{\text{bins}}=128$.
Hyperchaotic system simulations
As described above, the system was evolved according to [121]
$\dot{x}$ $=\alpha(y-x)+w$
$\dot{y}$ $=x(\beta-z)-y$
$\dot{z}$ $=xy-\gamma z$
$\dot{w}$ $=\delta w+\frac{1}{2}y^{2}-xz$
where $(\alpha,\beta,\gamma,\delta)=(10,28,2.7,2.2)$. We took $10^{2}$ initial conditions, which were allowed a burn-in time of $t_{\text{burn-in}}=1000$ steps to reach the attractor, after which they were simulated for $10^{5}$ steps with $\text{d}t=10^{-2}$.
Fluid flow simulations
The fluid flow simulations simulations are contained in the “Cylinder in Crossflow” dataset, downloaded from Ref. [142]. We take the dataset at Reynolds number 150, and interpolate the velocity field from the unstructured 6569-node mesh to a regular grid of $300\times 150$ pixels. The VIB networks are trained with $\beta=10^{7}$ and a time buffer $\Delta t$ chosen randomly between $[4,24]$ (see our comment on time randomization in “VIB for deterministic dynamics” in Section V).
Locally-coupled Kuramoto model
In SI Fig. S16 we perform variational IB on a locally-coupled Kuramoto model and find the latent variables learn a synchronization order parameter as in the cyanobacteria data. This model considers a spatially-distributed field of phases which evolves according to
@@
X Comparison of various dimensionality reduction methods
Variational IB (VIB) is by no means the only numerical method for performing data-driven model reduction. In this section we compare the performance of variational IB (VIB) to several common data-driven model reduction or inference methods. VIB uses a neural network to identify the relevant variables that are most predictive of the future. This construction learns dynamically relevant variables in contrast to methods such as principal component analysis or diffusion maps (in the absence of additional stringent assumptions), and it finds potentially non-linear variables in an agnostic way without requiring a tailored library of non-linearities (such as extended dynamic mode decomposition). In addition to these factors, VIB produces physically well-defined relevant variables – transfer operator eigenfunctions – in contrast to other deep methods such as (variational) autoencoders. In Fig. S17 we compare the low-dimensional latent space trajectories of the reduced models obtained using several of these standard methods to those output by VIB. In the remainder of this section we discuss these methods in detail.
@@ X Comparison of various dimensionality reduction methods
The principal components of the data are eigenvectors of the covariance matrix; the dominant principal component determines the direction of most variation in the dataset. In practice, if many samples are present the correlation matrix is expensive to compute. Writing the dataset as a matrix, $\mathbf{X}\in\mathbb{R}^{N_{\text{samples}}\times N_{\text{feat}}}$, computing the correlation matrix requires evaluating the matrix product $\mathbf{C}\propto\mathbf{X}^{T}\mathbf{X}$. Instead, a more computationally efficient way is to compute the singular value decomposition (SVD) of the data matrix,
@@
DMD in its original formulation assumes that the Koopman operator linearly evolves the observed state variable. However, the true Koopman operator evolves arbitrary non-linear functions of the state variable forward in time. DMD has therefore been extended to account for non-linear functions through “extended DMD”, or eDMD [17]. eDMD augments the state vector with non-linear transformations of the state. As an example, a state vector $\mathbf{x}-]$ may be replaced by monomials $\mathbf{x}-,-\mathbf{x}^{2}-,-\mathbf{x}^{3}-]$ (where here we understand exponentiation as element-wise). Other approaches also exist, such as augmenting the state vector by several time-delayed state vectors [144]. Rather than constructing the full $d\times d$ matrix, where $d$ is the dimension of the (possibly augmented) state, one can directly construct a low-rank approximation of $K$ by computing the reduced-rank singular value decomposition (SVD) of the matrix $K$ from the SVDs of $\mathbf{X}_{t}$ and $\mathbf{X}_{t+\Delta t}$.
@@
$\mathbf{x}=\mathbf{A}_{\text{ICA}}\mathbf{s}.$ (S112)
@@
The leveraging of non-Gaussianity in ICA is required due to the fact that, if using whitened data ($\mathbf{z}=\mathbf{V}\mathbf{x}$ so that $\mathbf{z}$ is zero mean $\mathbb{E}[\mathbf{z}]=\mathbf{0}$, and unit variance, $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=\mathbf{I}$), then the mixing matrix $\mathbf{A}_{\text{ICA}}$ cannot be estimated if the components $\mathbf{s}$ are normal Gaussian variables: any orthogonal matrix will satisfy (S112). For dynamical systems, however, one can get around this requirement of non-Gaussianity. In particular, one can use a time-lagged variable $\mathbf{z}_{t-\tau}=\mathbf{A}_{\text{ICA}}\mathbf{s}_{t-\tau}$ to compute the mixing matrix from the time-correlation matrix of the signal $\mathbf{z}_{t}$ [24]. In particular, we have
@@
where $q_{\phi}(z|x)$ is the posterior on the latent variables $z$ and is parameterized by a neural network (encoder) with parameters $\phi$, and $p_{\theta}(x|z)$ denotes the decoding network with parameters $\theta$. Strictly speaking, we present in the above objective function the $\beta$-VAE loss [112]. With $\beta=1$, which is the case for the original VAE [148], the objective is an upper bound on the log likelihood $\mathbb{E}[\log p_{\theta}(x)]$ which is minimized when $p_{\theta}(x)$ is equal to the true data distribution $p(x)$. The term $\beta$ controls compression as in the IB objective, however it has the opposite effect: small $\beta_{\text{IB}}$ corresponds to high compression, while small $\beta_{\text{VAE}}$ corresponds to low compression.